# A Survey on Monocular Re-Localization: From the Perspective of Scene Map Representation

### v1

@@ 4.1 Local Feature Extraction-then-Matching
Local Feature Extraction: As previously mentioned in Sec. 3.1, local feature detects massive salient pixels in an image (denoted as key points), and describes the neighboring area of the key point using a high-dimensional vector (denoted as descriptor). Under the context of VL-MRL methods, local feature is utilized to obtain accurate 2D-2D pixel-wise correspondences between images. The key points detected by local feature algorithms should be repeatable under different image conditions and qualities while located in salient regions zhang2017learning so that visual landmarks built by local features can be sparse but informative to represent the whole scene rapnet ; DSFeat . And descriptors should be designated to enable accurate correspondence establishment between detected feature points across different views l2net ; hardnet ; sosnet ; geodesc ; contextdesc , so they need be invariant against typical visual interference. Additionally, local feature should be accurate.
The early local features are built upon the human prior knowledge about the quality of pixels. As milestones in this area, the float SIFT sift , SURF surf and binary BRISK brisk , BRIEF brief , ORB orb have been widely used and achieved great successes in many autonomous applications. However, due to the limited presentation ability, these hand-crafted local features are challenged by complex visual interference in real-world scenarios. Since entering the perception age ruled by DNN, researchers have begun to introduce deep learning technique in local feature extraction. As a beginning, SuperPoint superpoint utilizes a shared encoder and two heads to automatically detect key points and extract descriptors, respectively. It is first trained on synthetic samples and then fine-tuned in two images generated by the homographic adaptation. After that, the training datasets with posed images and 3D map are constantly proposed, the ground-truth correspondence between different views of the scene can be available, facilitating the training of DNN-based local features. D2-Net d2net is optimize by apply metric learning between two co-visible patches and achieve better performance, but it extracts local feature on a feature map with reduced resolution, so the accuracy of matching and localization is limited. ASLFeat aslfeat improves by detecting on feature map with original resolution. ALIKE alike and ALIKED aliked support sub-pixel-level detection. Thus, their accuracy are boosted with a lot of margin. As a commonly applied strategy, multiple scale (MS) detection can help local feature extraction. Some works explicitly apply MS detection d2net ; sekd , some other works implicitly apply MS detection in network by fusing multiple features from feature pyramid inherited by CNN model aslfeat ; alike . To make DNN-based local feature get rid of costly training data collection, SEKD sekd proposes a self-evolving framework to train CNN model without any annotations or pre-processing on the training data. Some researchers apply attention mechanism in CNN model to further enhance the representative ability of local features delf ; mda . These methods train CNN model as GF-VPR tasks (see Sec. 3.1) and then detect local features based on estimated attention mask of trained CNN model. To facilitate the interpretability of trained attention mask, DSFeat DSFeat further combine attention information and semantic information to select stable and discriminative local features. Different from the methods trained with metric learning scheme, DISK disk train CNN model by reinforcement learning scheme where the feature extraction and matching procedures can be trained in an E2E manner, thereby greatly boosting the performance of local features.
@@
Figure 8: The processed Frame Per Second (FPS) and Mean Matching Accuracy (MMA) on HPatches dataset hpatches1 ; hpatches2 . The FPS metrics of SIFT sift , SURF surf , D2-Net d2net , SEKD sekd , GLFeat glfeat are evaluated on a NVIDIA RTX GPU glfeat , whereas LF-Net lfnet , SuperPoint superpoint , R2D2 r2d2 , ASLFeat aslfeat , DISK disk , ALIKE alike , and ALIKED aliked are evaluated on a NVIDIA GeForce RTX 2060 GPU aliked . “SS” denoted single scale detection while “MS” is multiple scale detection.
@@
Limited by the locality of CNN, most existing local features description methods only learn local descriptors with local information and lack awareness of global and surrounding spatial context. Thus, MTLDesc MTLDesc and GLFeat glfeat aggregate non-local awareness into local feature extraction. Besides, as shown in Fig. 8, most of the state-of-the-art local features cannot meet the requirements of efficiency even with high-performance GPU devices, which blocks the usage of local feature algorithms in real-time applications. Zhao et al. specially designed light CNN model for real-time local features alike ; aliked . Thin neural networks usually sacrifice representative ability, so strong supervision for local feature training is required to achieve balance between matching accuracy and computational efficiency alike ; aliked ; glfeat .
@@
Local Feature Matching: Before deep learning stage, researchers often apply the nearest neighbor (NN) searching to find matches between the extracted local features and perform ratio test to reject mismatches sift . This strategy is then improved by using the Earth Mover’s Distance Pele2008ECCV . Then, FLANN algorithm FLANN is developed as a efficient alternate of brute-force searching. To get more robust matches, various methods explores local key points distribution Guo2012PRL ; Ma2014TIP ; LPM ; Ma2018GLP ; CODE ; GMS . In addition to local key points distribution exploration, other strategies such as suitable local feature selection Hu2015TIP and optical flow guided matching Maier2016GuidedMB have also been explored to recover more matches. However, the above methods could ignore the global consistency since they only focus on the local distribution. On the other hand, the global properties of key points have also been explored ICF ; Yaron2014ToG ; RIRW ; DCM ; SOGM ; Lorenzo2008ECCV ; BBHomo . Moreover, the feature matching can be integrated into a RANSAC loop ransac to obtain matches with most inliers. Recently, AdaLAM adalam has been proposed to detects inliers by searching for significant local affine patterns in image correspondences. However, these traditional matching methods are less computational efficient due to hand-crafted rules and iterative formulation, and the performance will degrade if the underlying matching model between local feature in real scenarios differs from the pre-defined model of matching algorithms. In VPS , Sattler et al. proposed a novel vocabulary-based prioritized matching step that enables to first consider features more likely to yield 2D-to-3D matches and to terminate the correspondence search as soon as enough matches have been found, making VL-MRL method much more efficient and effective.
@@
As for learning-based local feature matching methods, Some algorithms explore semantic information to match the semantic key points in different instances of the same category of objects DeepSemanticFM ; Khan2020TCSVT ; Yu2018HierarchicalSemanticIM . Inspired by PointNetPointNet , a DNN of the point set learning, putative matches filtering networks good_correspondence ; NMNet ; OANet ; ACNe are developed. They commonly organize the putative matches of key points as 4D quads and feed the 4D quads into network to estimate scores of each putative match. False matches are discarded by a threshold on the scores. These methods merely consider the position information of the key point and neglect the descriptor of local features.
@@
In 2019, SuperGlue superglue was proposed as a E2E learning-based approach for local feature matching. SuperGlue superglue takes advantage of both key points and descriptors of local features and learns their matches with a graph neural network (GNN). The descriptor of each local feature is refined with all the features in both images by a attentional GNN layer so that matching will be reliable. Later, Zhao et al. IMUPrior exploited motion estimation from IMU integration and use the estimation as a spatial distribution prior of key points between images. With the assistance of such a prior, the network can make less efforts to refine features and achieve comparable accuracy with less GNN layers, boosting the real-time performance. More recently, LightGlue lindenberger2023lightglue improves the inference speed by enabling the model prune unmatchable points and predict whether further computation is required.
@@
4.2 Joint Local Feature Extraction and Matching
Technically similar to SIFT Flow SIFTFLow , joint local feature extraction-and-matching methods integrate feature extraction and matching module in an unified DNN model, and they directly estimate dense local feature matches so that the two module can be jointly optimized, as shown in Fig. 9 b). Early proposals in this area adopt convolutional DNN (CNN) model, e.g., NCNet NCNet , Sparse NCNet SparseNCNet , and DRC-Net DRCNet . Although all the possible matches should be considered in the CNN of these works, the receptive field of convolution is still limited to neighborhood area, making matching sided and unliable.
@@ 4.2 Joint Local Feature Extraction and Matching
To solve the problem, in recent three years, some works begin to adopt global attention mechanism, Transformer transformer , to achieve global consensus between matches. As the first proposal, COTR COTR applies zooming in method to obtain more accurate matching. LoFTR LoFTR proposes to perform dense matching in a coarse-to-fine framework, that it firstly obtains confidence matrices for patch-level coarse matching and then refines matched points positions at pixel level. Both steps using self- and cross-attention layers in Transformer transformer to enhance the representative ability of feature descriptors. It achieves accurate matching performance even in textureless region, as shown in Fig. 10, effectively reduce the requirements of scene texture in VL-MRL methods. As follow-up methods, QuadTree LoFTR Tang2022QuadTreeAF , ASpanFormer Chen2022ASpanFormerDI , and TKwinFormer Liao2023TKwinFormerTK make efforts to improve the matching accuracy and efficiency.
Figure 10: Some instances of local feature matches by SuperPoint superpoint +SuperGlue superglue , DRCNet DRCNet , and LoFTR LoFTR . The picture is from LoFTR .
@@
Existing matching methods choose to ignore the occluded objects caused by camera motion and scene structure, and they will fail when there is no overlap region between $\mathcal{I}_{q}$ and $\mathcal{I}_{r}$. To overcome this limitation, $\mbox{OCC}^{2}\mbox{Net}$ Fan2023Occ2NetRI models occlusion relations using 3D occupancy, enabling local feature matching for occluded regions.
@@
Summary: The joint extraction-and-matching scheme breaks through the limits that matching need to first extract excellent local features, and achieves impressive matching performance even in challenging scenarios, making VL-MRL methods more practical. But it need both $\mathcal{I}_{q}$ and $\mathcal{I}_{r}$ as inputs, which requires the scene map additionally saves $\mathcal{I}_{r}$, enlarging the storage cost of scene map. A storage-friendly scheme need to be explored.
4.3 Pose Solver
@@
Given the camera instrinsic $K$ and camera pose $\textbf{x}^{\mathcal{G}}_{q}$, the 2D camera projection $\textbf{p}^{\mathcal{I}_{q}}_{i}=\left[u_{i},v_{i}\right]$ from 3D visual landmarks $\textbf{m}_{i}=<\textbf{p}^{\mathcal{G}}_{i},*>,\textbf{p}^{\mathcal{G}}_{i}=\left[X^{\mathcal{G}}_{i},Y^{\mathcal{G}}_{i},Z^{\mathcal{G}}_{i}\right]$ to image plane $\mathcal{I}_{q}$ can be formatted as:
@@
$\begin{split}\left[\begin{array}[]{c}\textbf{p}^{\mathcal{I}_{q}}_{i}\\ 1\end{array}\right]&=\frac{1}{Z^{\mathcal{C}}_{i}}K\left[X^{\mathcal{C}}_{i}~{}Y^{\mathcal{C}}_{i}~{}Z^{\mathcal{C}}_{i}\right]^{T}\\ &=\frac{1}{Z^{\mathcal{C}}_{i}}K\left[\textbf{R}^{\mathcal{C}}_{\mathcal{G}}~{}|~{}\textbf{t}^{\mathcal{C}}_{\mathcal{G}}\right]\left[X^{\mathcal{G}}_{i}~{}Y^{\mathcal{G}}_{i}~{}Z^{\mathcal{G}}_{i}~{}1\right]^{T}\\ &=\frac{1}{Z^{\mathcal{C}}_{i}}K\left(\textbf{x}^{\mathcal{G}}_{\mathcal{C}}\right)^{-1}\left[\begin{array}[]{c}\textbf{p}^{\mathcal{G}}_{i}\\ 1\end{array}\right]\end{split}$ (12)
@@
where {}${\{\cdot\}}^{\mathcal{C}}$ denotes points in camera frame $\mathcal{C}$. And PnP problem is to estimate $\hat{\textbf{x}}^{\mathcal{G}}_{q}$ given $K$ and $n$ 2D-3D correspondences $\{\textbf{p}^{\mathcal{I}_{q}}_{i},\textbf{p}^{\mathcal{G}}_{i}\}^{n}_{i=1}$. This is a typical problem in 3D vision. Direct Linear Transformation (DLT) was first developed by photogrammetrists DLT to solve PnP problem and then introduced to computer vision community MVG ; 3DCV . If the 3D visual landmarks are coplanar, the homography transformation can be used instead MVG . In the last two decades there has been a great deal of progress in PnP solvers. The minimum number of correspondences to solve PnP prblem is reduced from 5 P5P , 4 P4P , to 3 p3p2 ; p3p1 . And the prominent solver is P3P algorithm ransac ; p3p2 ; p3p1 . Most of the solvers are able to deal with an arbitrary number of points. These generic PnP solvers are categorised as non-iterative and iterative.
Early non-iterative methods are generally computationally expensive, e.g., $O(n^{5})$ Quan1999TPAMI and $O(n^{2})$ Fiore2001TPAMI . The EPnP solution Lepetit2009EPnPAA is the first widely known $O(n)$ solver for the PnP problem. It represents 3D point coordinates as a linear combination of four control points. These control points are then retrieved using linearization techniques. Apart from linearization techniques, polynomial solvers have also been explored and become mainstream RPnP ; DLS ; OPnP ; SRPnP ; EOPnP . These works all estimate the camera pose by solving polynomial equations with $O(n)$ complexity.
Iterative methods are used to solve PnP optimization problems in an iterative manner, and it requires an appropriate initial values LHM ; FP ; SQPnP . The non-iterative solutions can be employed to provide an initial values, and iterative algorithms, such as Gauss-Newton (GN) iterations, are then used to refine the estimation. For instance, EPnP Lepetit2009EPnPAA and MLPnP Urban2016MLPnPA both enhance the performance by adding the GN iterations.
@@
The PnP problem can also be seen as a non-linear least square problems with regard to re-projection error so that camera pose can be solved by Bundle Adjustment (BA) orbslam3 . Looking back to Eqn. 12, the re-projection error can be formatted as:
$\begin{split}e_{i}=\overline{\textbf{p}}^{\mathcal{I}_{q}}_{i}-\frac{1}{Z^{\mathcal{C}}_{i}}K\left(\textbf{x}^{\mathcal{G}}_{\mathcal{C}}\right)^{-1}\overline{\textbf{p}}^{\mathcal{G}}_{i}\end{split}$ (13)
@@
where $\overline{\textbf{p}}$ means homogeneous coordinates of p. To estimate optimal camera pose, we need to minimize the total re-projection errors:
$\begin{split}\hat{\textbf{x}}^{\mathcal{G}}_{q}=\mathop{\arg\min}\limits_{\textbf{x}^{\mathcal{G}}_{q}}\frac{1}{2}\sum^{n}_{i=1}e^{2}_{i}\end{split}$ (14)
@@
And the function can be iteratively done by non-linear optimization methods like GN method and Levenberg-Marquardt (LM) method. In ORB-SLAM3 orbslam3 , the VL-MRL after tracking lost or kidnapping is implemented by first using MLPnP solver Urban2016MLPnPA to estimate an initial result and then refining with BA optimization.
@@
Summary: Given accurate 2D-3D correspondences, PnP is a well-established problem and can be excellently solved. Pose can be precisely estimated. However, in most real-world cases, 2D-3D correspondences usually have many or even predominant mismatches, challenging the pose estimation procedure. The PnP solver is always integrated into a RANSAC loops ransac to estimate more robustly and precisely hloc2 ; orbslam3 .
4.4 Further Improvements

### v3

@@ IV-A Local Feature Extraction-then-Matching
Local Feature Extraction: Local feature detects massive salient pixels in an image (denoted as key points), and describes the neighboring area of the key point using a high-dimensional vector (denoted as descriptor). Under the context of VL-MRL methods, local feature is utilized to obtain accurate 2D-2D pixel-wise correspondences between images. The key points detected by local feature algorithms should be repeatable under different image conditions and qualities while located in salient regions [97] so that visual landmarks built by local features can be sparse but informative to represent the whole scene [40, 68]. And descriptors should be designated to enable accurate correspondence establishment between detected local feature across different views [121, 122, 123, 124, 125], so they need be invariant against typical visual interference. Additionally, the positions of local features should be accurate [126, 95, 96].
The early local features are built upon the human prior knowledge about the quality of pixels. As milestones in this area, float-valued SIFT [56], SURF [70] and binary-valued BRISK [127], BRIEF [128], ORB [69] have been widely used and achieved great successes. However, due to the limited presentation ability, these hand-crafted local features are challenged by complex visual interference in real-world scenarios. Since entering the perception age ruled by DNN, researchers have begun to introduce DL technique in local feature extraction. As a milestone, SuperPoint [57] was proposed in 2018 and it leverage a Convolutional Neural Network (CNN) model to extract local features in an E2E way. The utilized CNN model is composed of a shared encoder and two heads. The encoder extracts high-dimensional dense features, and one head generate pixel-wise score map to detect key points while the other head generate descriptors. Such a network architecture is widely used in later works [97, 129, 95, 96, 94] and called “detect-then-describe” framework. D2-Net [130] and ASLFeat [126] adopt a different CNN architecture that only a encoder is used to generate dense descriptor map, and the score map used to detect key point is calculated from the descriptor map, called “detect-and-describe” framework.
The most important part in designing DL-based local feature is its training scheme, which determines the performance of local features, e.g., repeatability and accuracy of key points, robustness and discriminativeness of descriptors. SuperPoint [57], as a ganger, is first trained on synthetic samples and then fine-tuned in two images generated by its proposed homographic adaptation techinique. Later, some real-world datasets [131] with posed images and 3D map are proposed, the ground-truth correspondences between different images can be available, facilitating the training of DL-based local features. Most algorithms apply metric learning between two co-visible patches or images to optimize their network parameters so that the network can gradually learn to extract local features suitable for feature matching [130, 126, 95, 96, 97]. So, their key points are repeatable while descriptors can be precisely matched. To make DNN-based local feature get rid of costly training data collection, SEKD [129] proposes a self-evolving framework to train CNN model without any annotations or pre-processing on the training data. DISK [94] trains CNN model not by metric learning but by reinforcement learning scheme where the feature extraction and matching procedures can be trained in an E2E manner, thereby greatly boosting the performance of local features. To detect informative key points in images, some methods regard the score map as attention mask to aggregate dense descriptor, so the local feature algorithms can be trained as GF-VPR methods, which also brings significant improvements [132, 133, 68].
@@
Figure 8: The processed Frame Per Second (FPS) and Mean Matching Accuracy (MMA) on HPatches dataset [134, 135]. The FPS metrics of SIFT [56], SURF [70], D2-Net [130], SEKD [129], GLFeat [136] are evaluated on a NVIDIA RTX GPU [136], whereas LF-Net [137], SuperPoint [57], R2D2 [97], ASLFeat [126], DISK [94], ALIKE [95], and ALIKED [96] are evaluated on a NVIDIA GeForce RTX 2060 GPU [96]. “SS” denoted single scale detection while “MS” is multiple scale detection.
@@
Limited by the inherent locality of CNN, most existing local features description methods only learn local descriptors with local information and lack awareness of global context. Thus, MTLDesc [138] and GLFeat [136] aggregate non-local information into local feature extraction. Besides, as shown in Fig. 8, most of the State-of-the-Art (SOTA) local features cannot meet the requirements of efficiency even with high-performance GPU devices, which blocks the usage of local feature algorithms in real-time applications. Zhao et al. specially designed light CNN model for real-time local features [95, 96]. Thin neural networks usually sacrifice representative ability, so strong supervision for local feature training is required to achieve balance between matching accuracy and computational efficiency [95, 96, 136].
Monocular images lack of actual scale information, which affect the performance of local features. As a commonly applied strategy, Multiple Scale (MS) detection can help local feature extraction. Some works explicitly apply MS detection [130, 129], while some other works implicitly apply MS detection in network by fusing multiple features from feature pyramid inherited by CNN model [126, 95].
@@
Local Feature Matching: Before DL stage, researchers often apply the brute-force Nearest Neighbour (NN) searching to find matches between the extracted local features and perform ratio test to reject mismatches [56]. Then, FLANN algorithm [139] is developed as an efficient alternate of brute-force searching. To get more robust matches, some methods explore positional distribution of local features [140, 141, 142], and some other works further explore locally adaptive local descriptor selection [143] and optical flow guided matching [144] for better matching. These methods only consider local properties of local features so they can hardly achieve globally consistent matching. So, some approaches turn to leverage the global properties of local features [145, 146, 147, 148]. In most practical applications, a simple yet effective method is perform epipolar check after NN matching that estimate essential matrix between two images based on NN matched local features, and reject mismatched features being contrary to the underlying epipolar constraints formatted as essential matrix [149, 150, 120]. But if mismatched local features were in the majority, the estimated essential matrix will be wrong and this solution will fail. Thus, we usually integrated this solution into a RANSAC loop [118] to obtain estimated model with most inliers and use the model to reject mismatches. Recently, AdaLAM [99] has been proposed to detects inliers by searching for significant local affine patterns in image correspondences. These traditional hand-crafted matching methods have large running latency and require costly computational consumption due to complex rules and iterative formulation, and the performance will degrade if the underlying model between local features in real scenarios differs from the pre-defined model. In [151], Sattler et al. proposed a novel vocabulary-based prioritized matching step that enables to first consider features more likely to yield 2D-3D matches and to stop searching matches as soon as enough matches have been found, making local feature matching much more efficient and effective.
@@
As for DL-based local feature matching methods, some algorithms explore semantic information to match the semantic key points in different instances of the same category of objects [152, 153]. Some works regard local feature matching as a point set learning problem and develop putative matches filtering networks [102, 154, 155, 106]. They commonly organize the putative matches of key points as 4D quads and feed the 4D quads into network to estimate scores of each putative match. False matches are discarded by a threshold on the scores. These methods only consider the position information of the key points and neglect the descriptor information of local features, thus remaining room for improvements.
@@
In 2019, SuperGlue [98] was proposed as an E2E learning-based approach for local feature matching and achieved impressive success. SuperGlue [98] considers both key points and descriptors of local features and learns their matches with a Graph Neural Network (GNN). Each local features are modelled as a node in the graph and the descriptors are enhanced by all the features in both images by an attentional GNN layer so that matching will be reliable. Later, Zhao et al. [156] exploited motion estimation from IMU integration and use the estimation as a spatial distribution prior of matched local features between images. With the assistance of such a prior, the network can achieve comparable accuracy with less GNN layers, boosting the real-time performance. More recently, LightGlue [100] improves the inference speed by enabling the model prune unmatchable local features and predict whether further computation is required.
@@
IV-B Joint Local Feature Extraction and Matching
Joint local feature extraction-and-matching methods integrate feature extraction and matching module in a unified DNN model, and they directly estimate dense local feature matches so that the two modules can be jointly optimized in an E2E way, as shown in Fig. 9 b). Early proposals in this area adopt CNN model, e.g., NCNet [157], Sparse NCNet [158], and DRC-Net [159]. Although all the possible matches should be considered in the CNN of these works, the receptive field of convolution is still limited to neighborhood area, making matching sided and unreliable.
@@ IV-B Joint Local Feature Extraction and Matching
To solve the problem, in recent three years, some works begin to adopt global attention mechanism by Transformer [65] to achieve global consensus between matches. As the first proposal in this line, COTR [160] use a Transformer-based model that queries a local feature in one image and estimates the matched local feature on another image. For efficiency, COTR [160] apply the network by recursively zooming in around the estimations. LoFTR [161] proposes to perform dense matching in a coarse-to-fine framework, that it firstly obtains confidence matrices for patch-level coarse matching and then refines matched points positions at pixel level. Both steps using self- and cross-attention layers by Transformer [65] to enhance the representative ability of feature descriptors. It achieves accurate matching performance even in textureless region, effectively reducing the requirements of scene texture in VL-MRL methods. As follow-up methods, some methods make efforts to improve the matching accuracy and efficiency [162, 163, 164].
@@
Existing matching methods choose to ignore the occluded area, and they will fail when there is no overlap region between $\mathcal{I}_{q}$ and $\mathcal{I}_{r}$. To overcome this limitation, $\mbox{OCC}^{2}\mbox{Net}$ [165] models occlusion relations using 3D occupancy, enabling local feature matching for occluded regions.
@@
Summary: The joint extraction-and-matching scheme breaks through the limits of extraction-then-matching schemes that good matching relies on excellent local features, and achieves impressive performance, making VL-MRL methods more accurate. But it requires the scene map additionally saves $\mathcal{I}_{r}$, enlarging the storage requirements of scene map.
IV-C Pose Solver
@@
Given the camera instrinsic $K$ and camera pose $\textbf{x}^{\mathcal{G}}_{q}$, the 2D camera projection $\textbf{p}^{q}_{i}=\left[u,v\right]$ from 3D visual landmarks $\textbf{m}_{i}=<\textbf{P}_{i},*>,\textbf{P}_{i}=\left[X^{\mathcal{G}}_{i},Y^{\mathcal{G}}_{i},Z^{\mathcal{G}}_{i}\right]$ to image plane $\mathcal{I}_{q}$ can be formatted as:
@@
$\begin{split}\left[\begin{array}[]{c}\textbf{p}^{q}_{i}\\ 1\end{array}\right]&=\frac{1}{Z^{\mathcal{C}}_{i}}K\left[X^{\mathcal{C}}_{i}~{}Y^{\mathcal{C}}_{i}~{}Z^{\mathcal{C}}_{i}\right]^{T}\\ &=\frac{1}{Z^{\mathcal{C}}_{i}}K\left[\textbf{R}^{\mathcal{C}}_{\mathcal{G}}~{}|~{}\textbf{t}^{\mathcal{C}}_{\mathcal{G}}\right]\left[X^{\mathcal{G}}_{i}~{}Y^{\mathcal{G}}_{i}~{}Z^{\mathcal{G}}_{i}\right]^{T}\\ &=\frac{1}{Z^{\mathcal{C}}_{i}}K\left(\textbf{x}^{\mathcal{G}}_{\mathcal{C}}\right)^{-1}\textbf{P}_{i}\end{split}$ (16)
@@
where {}${\{\cdot\}}^{\mathcal{C}}$ denotes points in camera coordinate $\mathcal{C}$. And PnP problem is to estimate $\hat{\textbf{x}}^{\mathcal{G}}_{q}$ given $K$ and $n$ 2D-3D correspondences $\{\textbf{p}^{q}_{i},\textbf{P}_{i}\}^{n}_{i=1}$. This is a typical problem in 3D vision. Direct Linear Transformation (DLT) [166] was first developed by photogrammetrists to solve PnP problem and then introduced to computer vision community. In the last two decades, PnP solvers have been well studied. The minimum number of correspondences to solve PnP problem is reduced from 5 [167], 4 [168], to 3 [169, 170].
Generally, the PnP solvers can be categorized as non-iterative and iterative. Early non-iterative PnP solvers are generally computationally expensive, e.g., $O(n^{5})$ [171]. The EPnP solution [172] effectively reduces the computation complexity to $O(n)$ by representing 3D point coordinates as a linear combination of four control points, which is widely utilized in many MRL solutions. In addition to EPnP [172], some other non-iterative PnP solvers with $O(n)$ complexity are later proposed [173, 174, 175, 176] but they are polynomial solvers compared to linearization-based EPnP [172]. Instead, iterative methods solve PnP problems in an iterative manner, beginning from an appropriate initial value [177, 178, 179]. The non-iterative solutions are employed to provide an initial value, and iterative algorithms are used to refine the estimation [172].
@@
The PnP problem can also be seen as a non-linear least square problems with regard to re-projection error so that camera pose can be solved by Bundle Adjustment (BA) [120]. From Eqn. 16, the re-projection error is defined as:
$\begin{split}e_{i}=\textbf{p}^{q}_{i}-(\frac{1}{Z^{\mathcal{C}}_{i}}K\left(\textbf{x}^{\mathcal{G}}_{\mathcal{C}}\right)^{-1}\textbf{P}_{i})[:2]\end{split}$ (17)
@@
To estimate optimal camera pose, we need to minimize the total re-projection errors:
$\begin{split}\hat{\textbf{x}}^{\mathcal{G}}_{q}=\hat{\textbf{x}}^{\mathcal{G}}_{\mathcal{C}}=\mathop{\arg\min}\limits_{\textbf{x}^{\mathcal{G}}_{\mathcal{C}}}\frac{1}{2}\sum^{n}_{i=1}e^{2}_{i}\end{split}$ (18)
@@
And the function can be iteratively done by non-linear optimization methods like Gauss-Newton (GN) method and Levenberg-Marquardt (LM) method. In ORB-SLAM3 [120], the VL-MRL after tracking lost or kidnapping is implemented by first using MLPnP solver [180] to estimate an initial result and then refining with BA optimization.
@@
Summary: Given accurate 2D-3D correspondences, PnP is a well-established problem and can be excellently solved. Pose can be precisely solved. However, in most real-world cases, 2D-3D correspondences usually have many or even predominant mismatches, challenging the pose estimation procedure. The PnP solver is always integrated into a RANSAC loops [118] to estimate more robustly and precisely [24, 120].
IV-D Further Improvements